for a bakery the cost of making each day's batch of doughnuts is given by the equation below, where C is the total cost in dollars of making a batch of doughnuts and N is the number of doughnuts in the batch.
C=30+0.1N
If the number of doughnuts in a batch is increased by 20 dougnuts, by how much will the cost of the batch increse?

Answer by rajagopalan(158)

cost of making each day's batch of N doughnuts=C=30+0.1N
cost of making each day's batch of N+20 doughnuts=C=30+0.1(N+20)
Increase in cost ( step2-step1=30+0.1(N+20) minus 30+0.1N
=30+0.1N+(0.1*20)-30-0.1*N
=30+0.1N+2-30-0.1N
30and -30 cancel
0.1N and -0.1N cancel
Leaving 2 as remainder
So Additional cost =2 Dollars
The batch cost will increase by only 2 dollars
okay